OpenStudy (anonymous):

Number theory: "Use Wilson's theorem (p-1)! congruent to -1(mod p) to show that (p-3)1 congruent to ((p-1)/2) mod p" Thanks :)

OpenStudy (amistre64):

the only thing that comes to mind at the moment is tom hanks yelling at a volleyball :/

OpenStudy (amistre64):

i spose Wilsons thrm is: (p-1)! = -1 (mod p)

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

-3 = -1-2

OpenStudy (amistre64):

([p-2] -1) ! = -1 (mod [p-2]) but that doesnt seem right to me .. just yet

OpenStudy (amistre64):

(p-1)! = (p-1)(p-2)(p-3)!

OpenStudy (amistre64):

that should prolly be used to start out with

OpenStudy (amistre64):

multiply both sides by p-1 p-2, or i might be working it backwards ..

OpenStudy (amistre64):

(p-1)! = -1(mod p) (p-1)(p-2)(p-3)! = -1 (mod p) need to develop some inverses

OpenStudy (amistre64):

1 2 3 4 5 1 1 2 3 4 0 2 2 4 1 3 0 3 3 1 4 2 0 4 4 3 2 1 0 5 0 0 0 0 0 seems to be a good representation of a multiplication table in mod 5

OpenStudy (amistre64):

4! = -1 (m5) 432! = -1 (m5) 4432! = -14 (m5) 232! = -142 (m5) 2! = -8 (m5) -8+5+5 = 2 2! = 2 (m5) = (4-1)/2

OpenStudy (amistre64):

(p-1)(p-1)/p p -2 ----------- p | p^2-2p+1 (p^2) ------------ -2p+ 1 (-2p) -------- 1 <-- remainder of 1 (p-1)(p-1) = 1 (mod p)

OpenStudy (amistre64):

(p-2)(p-k)/p p -(2+k) ------------------ p | p^2 -(2+k)p +2k (p^2) ----------------- -(2+k)p +2k [(-2+k)p] -------------- 2k (mod p) for some k not too sure if that can help us out .....

OpenStudy (amistre64):

(p-1)(p-2)(p-3)! = -1 (mod p) ; lets use -1 = (p-1) (p-1)(p-2)(p-3)! = (p-1) (mod p) (p-1)(p-1)(p-2)(p-3)! = (p-1)(p-1) (mod p) (p-2)(p-3)! = 1 (mod p) (p-k)(p-2)(p-3)! = (p-k)1 (mod p) (p-3)! = (p-k) (mod p) (p-3)! = 2k (mod p)

OpenStudy (amistre64):

thats wrong a little bit i think ... but is it giving you any ideas?

OpenStudy (amistre64):

(p-k) = k (mod p)

OpenStudy (amistre64):

if we can show that k = (p-1)/2 youd be set

OpenStudy (amistre64):

all thats really left to show .. is that \[\Large \frac{(p-1)}{2}(p-2)\equiv1~(mod ~p)\] (p-1)(p-2) = p^2 -3p +2 p - 3 --------------- p | p^2 -3p +2 (p^2) ------ -3p (-3p) ------ 2; and 1/2 * 2 = 1 for a remainder

OpenStudy (anonymous):

I get it! Thankyou so much! I think it was the factorial sign that was throwing me off!

OpenStudy (amistre64):

it was a little something else throwing me at first :) but i figured it out